### v1

# Hydrostatic Newton-Cartan Membranes

Abstract
We construct the most general equilibrium partition function for Newton-Cartan membranes of codimension-1 defined on a flat ambient Newton-Cartan spacetime at second order in the hydrodynamic derivative expansion. We find the equilibrium constraints and the corresponding set of equilibrium constitutive relations. We show the explicit form of the Young-Laplace equation for a subset of non-vanishing and non-constant transport coefficients in the presence of a rotating fluid background with a constant vorticity.
I Introduction
Surfaces and interfaces are of importance on a wide range of physical phenomena such as biological systems Canham1970 ; Helfrich ; Tu2014 ; Cowin:2001 ; Mooendarbary:2013 ; Mow:1984 , condensed matter systems Kane2005 , and wave dynamics on spherical surfaces Streichan2017 ; Ritsma2014 ; Delplace2017 ; Shankar2017 ; Pearce2019 , such as planet Earth. A theoretical framework for studying such systems was established in Armas:2020 , based on the microscopic symmetry structure of the systems under study. A key aspect of the formalism used in Armas:2020 is that a geometric interpretation of temperature and mass chemical potential arise naturally from it. This allows for the study of elastic surfaces at finite temperature and mass chemical potential.
Following the work laid down in Armas:2020 , it is the goal of this paper to establish the equilibrium states of membranes that allow for exchange of both energy and matter. In particular, we will use a generating equilibrium partition function to write down the most general set of constitutive relations for membranes with finite temperature and chemical potential.
The manuscript is organized as follows. In sections II and III we provide a summary of the necessary tools to understand the formalism presented in Armas:2020 . Section IV shows the use of the formalism to write down the equilibrium partition function up to second order in derivatives together with its corresponding equilibrium constraints and constitutive relations. In section V we write down the corresponding Young-Laplace equation in the presence of constant vorticity for a particular set of non-vanishing transport coefficient that depend on both temperature and chemical potential. Finally, in sectionVI we conclude with some comments about potential future work.
II Torsional Newton Cartan Background
In this section we present an overview of Torsional Newton Cartan (TNC) geometry and introduce all the essential details that will be needed when using this geometry as the source of the mass, momentum, and energy currents. These details include a summary of TNC sub-manifolds, formalism that is needed to describe the equilibrium state of non-relativistic membranes. See Armas:2020 for a review [^1] Most of the overview presented in this and the next section, particularly regarding TNC sub-manifolds, can be considered a summary of what is discussed and introduced in Armas:2020 . of TNC geometry and TNC sub-manifolds.
II.1 TNC Geometry
Let $\mathcal{M}_{d+1}$ be a (d+1)-dimensional manifold equipped with a TNC structure, consisting on the fields [^2] We use Greek indices to denote space-time indices $\{\mu,\nu,...\}=0,...,d$ $\left(\tau_{\mu},h_{\mu\nu},m_{\mu}\right)$; with $h_{\mu\nu}$ a symmetric spatial metric, $\tau_{\mu}$ a 1-form known as the clock 1-form, and $m_{\mu}$ a U(1) gauge connection associated to particle number conservation. An inverse structure $\left(\upsilon^{\mu},h^{\mu\nu}\right)$ can be defined following the orthonormality conditions
$\tau_{\mu}h^{\mu\nu}$ $=0,\qquad\qquad$ $\tau_{\mu}\upsilon^{\mu}$ $=-1,$ (1)
$h_{\mu\nu}\upsilon^{\nu}$ $=0,\qquad\qquad$ $h_{\mu\rho}h^{\rho\nu}-\tau_{\mu}\upsilon^{\nu}$ $=\delta^{\nu}_{\mu}.$ (2)
The TNC structure on $\mathcal{M}_{d+1}$ in terms of the fields $\left(\tau_{\mu},h_{\mu\nu},m_{\mu}\right)$ transform under diffeomorphisms (local coordinate transformations parametrized by $\xi^{\mu}$), U(1) (mass gauge transformations parametrized by $\sigma$), Milne boosts (local rotations and local Galilean boosts parametrized by $\lambda_{\mu}$) as
$\delta\tau_{\mu}$ $=\mathcal{L}_{\xi}\tau_{\mu}\,,$ (3)
$\delta h_{\mu\nu}$ $=\mathcal{L}_{\xi}h_{\mu\nu}+\lambda_{\mu}\tau_{\nu}+\lambda_{\nu}\tau_{\mu}\,,$ (4)
$\delta m_{\mu}$ $=\mathcal{L}_{\xi}m_{\mu}+\lambda_{\mu}+\partial_{\mu}\sigma.$ (5)
A spacetime measure $e$ can be constructed from this objects through the identification
$e=\sqrt{\det\left(h_{\mu\nu}-\tau_{\mu}\tau_{\nu}\right)}.$ (6)
It is sometimes more convenient to work with the explicitly Milne boost invariant structure $\{\tau_{\mu},\bar{h}_{\mu\nu},\Phi\}$ and inverse structure $\{\hat{\upsilon}^{\mu},h^{\mu\nu}\}$, defined as
$\bar{h}_{\mu\nu}$ $\equiv h_{\mu\nu}-\tau_{\mu}m_{\nu}-\tau_{\nu}m_{\mu}\,,$ (7)
$\hat{\upsilon}^{\mu}$ $\equiv\upsilon^{\mu}-h^{\mu\nu}m_{\nu}\,,$ (8)
$\Phi$ $\equiv-\hat{\upsilon}^{\mu}m_{\mu}+\frac{1}{2}h^{\mu\nu}m_{\mu}m_{\nu}\,,$ (9)
and satisfying the orthonormality conditions
$\tau_{\mu}h^{\mu\nu}$ $=0\qquad\qquad$ $\tau_{\mu}\hat{\upsilon}^{\mu}=-1$ (10)
$\bar{h}_{\mu\nu}\hat{\upsilon}^{\nu}$ $=2\Phi\tau_{\mu}\qquad\qquad$ $\bar{h}_{\mu\nu}h^{\nu\rho}-\tau_{\mu}\hat{\upsilon}^{\rho}=\delta^{\rho}_{\mu}\ ,$ (11)
and transforming under local U(1) mass transformation as
$\delta\bar{h}_{\mu\nu}$ $=-\tau_{\mu}\partial_{\nu}\sigma-\tau_{\nu}\partial_{\mu}\sigma\,,$ (12)
$\delta\hat{\upsilon}^{\mu}$ $=-h^{\mu\nu}\partial_{\nu}\sigma\,,$ (13)
$\delta\Phi$ $=-\hat{\upsilon}^{\mu}\partial_{\mu}\sigma.$ (14)
The rest of the structures transform only under diffeomorphisms in the usual way. Spacetime measure $e$ can be written in terms of these variables as
$e=\sqrt{\frac{\det\bar{h}}{2\Phi}}.$ (15)
To study dynamic evolution of systems coupled to a TNC geometry covariantly it is necessary to introduce a covariant derivative adapted to the generic curved TNC background. There exist no unique metric-compatible connection so we ask for compatibility with $\{\tau_{\mu},h^{\mu\nu}\}$, namely
$\nabla_{\mu}\tau_{\nu}=0\ \qquad\text{and}\qquad\nabla_{\mu}h^{\rho\sigma}=0\,,$ (16)
with $\nabla$ the covariant derivative constructed using the affine connection $\Gamma^{\rho}{}_{\mu\nu}$. A connection satisfying (16) and explicitely boost invariant is given by
$\Gamma^{\rho}{}_{\mu\nu}=-\hat{\upsilon}^{\rho}\partial_{\mu}\tau_{\nu}+\frac{1}{2}h^{\rho\sigma}\left(\partial_{\mu}\bar{h}_{\nu\sigma}+\partial_{\nu}\bar{h}_{\mu\sigma}-\partial_{\sigma}\bar{h}_{\mu\nu}\right)\,.$ (17)
Unlike the Riemannian Christofell connection, $\Gamma^{\rho}{}_{\mu\nu}$ as defined in (17) is generically torsionful
$\Gamma^{\lambda}{}_{[\mu\nu]}=-\frac{1}{2}\hat{\upsilon}^{\lambda}F_{\mu\nu}\,,$ (18)
where we defined the torsion two form $F_{\mu\nu}$ as
$F_{\mu\nu}=\partial_{\mu}\tau_{\nu}-\partial_{\nu}\tau_{\mu}.$ (19)
For most cases of physical interest such torsion vanishes [^3] This statement is equivalent to saying that for torsionless systems there exists a universal time.. The connection (17) is invariant under boosts, however it transforms under U(1) mass transformations as
$\delta\Gamma^{\rho}_{\mu\nu}=\frac{1}{2}h^{\rho\lambda}\left(F_{\mu\nu}\partial_{\lambda}\sigma+F_{\lambda\nu}\partial_{\mu}\sigma+F_{\lambda\nu}\partial_{\nu}\sigma\right).$ (20)
For a vanishing torsion the connection is invariant under both boost and U(1) transformations. The Riemann tensor $R^{\rho}{}_{\sigma\mu\nu}$ of the TNC background is defined in the standard way
$R^{\rho}{}_{\sigma\mu\nu}=\partial_{\mu}\Gamma^{\rho}_{\nu\sigma}-\partial_{\nu}\Gamma^{\rho}_{\mu\sigma}+\Gamma^{\rho}_{\mu\lambda}\Gamma^{\lambda}_{\nu\sigma}-\Gamma^{\rho}_{\nu\lambda}\Gamma^{\lambda}_{\mu\sigma}.$ (21)
II.2 TNC Submanifolds
A (p+1)-dimensional TNC submanifold $\Sigma_{p+1}$, with coordinates [^4] We will use the lower case latin indices $\{a,b,...=0,...,p\}$ to denote submanifold spacetime indices $\sigma^{a}$, embedded on a (d+1)-dimensional TNC manifold $\mathcal{M}_{d+1}$, with coordinates $x^{\mu}$, is parametrized by the map $X^{\mu}:\Sigma\rightarrow\mathcal{M},\mu=0,...,d$ relating the coordinates $\sigma^{a}$ on $\Sigma_{p+1}$ to the location $X^{\mu}(\sigma^{a})$ on $\mathcal{M}$. In other words, the embedding specifies the location of the surface on the ambient space $\mathcal{M}$.
Tangent vectors $u^{\mu}_{a}$ to the surface are defined from the embedding map as
$u^{\mu}_{a}$ $=\partial_{a}X^{\mu}\,.$ (22)
The normal 1-forms [^5] The upper case latin indices $\{I,J,...=1,...,d-p\}$ correspond to indices on the transverse directions to the surface. We will consider codimension 1 surfaces meaning there will only be a single transverse direction and no upper latin indices will be used beyond this section. $n^{I}_{\mu}$ will be implicitly defined through the relations
$n^{I}_{\mu}u^{\mu}_{a}$ $=0\,,$ (23)
$h^{\mu\nu}n^{I}_{\mu}n^{J}_{\nu}$ $=\delta^{IJ},$ (24)
where $I=1,...,d-p$. Notice that $n^{I}_{\mu}$ is not uniquely fixed by (23) as such relation is invariant under SO(d-p) transformations. The inverse tangent vector $u^{a}_{\mu}$ and inverse normal 1-form $n^{\mu}_{I}$ can be introduce through the orthonormality conditions
$\delta^{\mu}_{\nu}$ $=u^{\mu}_{a}u^{a}_{\nu}+n^{I}_{\nu}n^{\mu}_{I}\,,$ (25)
$u^{a}_{\mu}n^{\mu}_{I}$ $=0\,,$ (26)
$u^{\mu}_{a}u^{b}_{\mu}$ $=\delta^{b}_{a}\,,$ (27)
$n^{\mu}_{I}n^{J}_{\mu}$ $=\delta^{J}_{I}.$ (28)
On time-like submanifolds the normal vector $n^{\mu}_{I}$ satisfies $n^{\mu}_{I}\tau_{\mu}=0$, therefore we can explicitly set $n^{\mu I}=h^{\mu\nu}n^{I}_{\nu}$. We can use the tangent vectors and normal 1-forms, as well as their inverses, to project any tensor tangentially or orthogonally to the surface. In particular the relevant boost-invariant TNC structure induced on the surface will be given by the set of objects $\{\tau_{a},\bar{h}_{ab},\check{\Phi}\}$ and $\{\hat{\upsilon}^{a},h^{ab}\}$ related to the ambient TNC structures through the relations
$\tau_{a}$ $=u^{\mu}_{a}\tau_{\mu}\,,$ (29)
$\bar{h}_{ab}$ $=u^{\mu}_{a}u^{\nu}_{b}\bar{h}_{\mu\nu}\,,$ (30)
$\hat{\upsilon}^{a}$ $=u^{a}_{\mu}\hat{\upsilon}^{\mu}\,,$ (31)
$h^{ab}$ $=u^{a}_{\mu}u^{b}_{\nu}h^{\mu\nu}\,,$ (32)
$\check{\Phi}$ $=\Phi-\frac{1}{2}\hat{\upsilon}^{\mu}\hat{\upsilon}^{\nu}n^{I}_{\mu}n^{J}_{\nu}\delta_{IJ}\,,$ (33)
and subject to the orthonormality conditions
$\tau_{a}h^{ab}$ $=0,\qquad\qquad$ $\tau_{a}\hat{\upsilon}^{a}=-1,$ (34)
$\bar{h}_{ab}\hat{\upsilon}^{b}$ $=2\check{\Phi}\tau_{a},\qquad\qquad$ $\bar{h}_{ab}h^{bc}-\tau_{a}\hat{\upsilon}^{c}=\delta^{c}_{a}.$ (35)
This induced TNC structure is invariant under local Milne boosts while transforming under submanifold diffeomorphisms ($\zeta^{a}$), and U(1) gauge transformations $\sigma$ as
$\delta\tau_{a}$ $=\mathcal{L}_{\zeta}\tau_{a}\,,$ (36)
$\delta\bar{h}_{ab}$ $=\mathcal{L}_{\zeta}\bar{h}_{ab}-\tau_{a}\partial_{b}\sigma-\tau_{b}\partial_{a}\sigma\,,$ (37)
$\delta\hat{\upsilon}^{a}$ $=\mathcal{L}_{\zeta}\upsilon^{a}-h^{ab}\partial_{b}\sigma\,,$ (38)
$\delta h^{ab}$ $=\mathcal{L}_{\zeta}h^{ab}\,,$ (39)
$\delta\check{\Phi}$ $=\mathcal{L}_{\zeta}\check{\Phi}-\hat{\upsilon}^{a}\partial_{a}\sigma.$ (40)
Note that the underlying U(1) transformation is the one associated to the gauge connection $\check{m}_{a}$, $\delta\check{m}_{a}=\partial_{a}\sigma$ , related to the bulk one via
$\check{m}_{a}$ $=u^{\mu}_{a}m_{\mu}-\frac{1}{2}\upsilon^{I}\upsilon_{I}\tau_{a}.$ (41)
Covariant differentiation can be defined along the surface direction through the introduction of the covariant derivative $D_{a}$ satisfying the compatibility conditions
$D_{a}\tau_{b}$ $=0,\qquad\qquad$ $D_{a}h^{bc}$ $=0,$ (42)
$D_{a}\tau_{\mu}$ $=0,\qquad\qquad$ $D_{a}h^{\mu\nu}$ $=0.$ (43)
This covariant derivative will act on a mixed index tensor $T^{b\mu}$ as
$D_{a}T^{b\mu}=\partial_{a}T^{b\mu}+\gamma^{b}_{ac}T^{c\mu}+u^{\rho}_{a}\Gamma^{\mu}_{\rho\lambda}T^{b\lambda},$ (44)
with the surface affine connection given by
$\gamma^{c}_{ab}$ $=-\hat{\upsilon}^{c}\partial_{a}\tau_{b}+\frac{1}{2}h^{cd}\left(\partial_{a}\bar{h}_{bd}+\partial_{b}\bar{h}_{ad}-\partial_{d}\bar{h}_{ab}\right)$ (45)
$=\Gamma^{c}_{ab}+u^{c}_{\mu}\partial_{a}u^{\mu}_{b}=u^{c}_{\mu}u^{\nu}_{a}\nabla_{\nu}u^{\mu}_{b}\,,$ (46)
and associated torsion tensor $F_{ab}$ defined through
$2\gamma^{c}_{[ab]}=-\hat{\upsilon}^{c}F_{ab}.$ (47)
A submanifold Riemann tensor $\mathcal{R}^{a}{}_{bcd}$ can be constructed from $\gamma^{a}_{ab}$ through the standard form
$\mathcal{R}^{a}{}_{bcd}=\partial_{c}\gamma^{a}_{db}-\partial_{d}\gamma^{a}_{cb}+\gamma^{a}_{cf}\gamma^{f}_{db}-\gamma^{a}_{df}\gamma^{f}_{cb}.$ (48)
The boost invariant extrinsic curvature of the submanifold $K_{ab}{}^{I}$ can be defined from the covariant derivative of the tangential vector through
$K_{ab}{}^{I}=-u^{\mu}_{a}u^{\nu}_{b}\nabla_{(\mu}n^{I}_{\nu)}=n^{I}_{\mu}D_{a}u^{\mu}_{b}+\frac{1}{2}\hat{\upsilon}^{I}\tau_{ab},$ (49)
and transforming under mass U(1) transformations as
$\delta K_{ab}{}^{I}=\frac{1}{2}F_{Ia}\partial_{b}\sigma+\frac{1}{2}F_{Ib}\partial_{a}\sigma.$ (50)
Notice that the covariant derivative $D$ does not act on normal indices $I$. It is also possible to define a covariant derivative $\mathcal{D}$ that acts on all types of indices. In particular we will find that
$\mathcal{D}_{a}T^{I}=D_{a}T^{I}-\omega_{a}{}^{I}{}_{J}T^{J}.$ (51)
The tensor $\omega_{a}{}^{I}{}_{J}$ is called the external rotation tensor and can be interpreted as a SO(d-p) connection, it is defined from the covariant derivative of the normal 1-form
$\omega_{a}{}^{I}{}_{J}=n^{\mu}_{J}D_{a}n^{I}_{\mu}\,,$ (52)
it is antisymmetric in $I,J$ indices, and transforms under U(1) mass transformations as
$\delta\omega_{a}{}^{I}_{J}=-\frac{1}{2}\left(F_{aJ}\partial^{J}\sigma+F^{I}{}_{J}\partial_{a}\sigma+\tau^{I}{}_{a}\partial_{J}\sigma\right).$ (53)
Notice that for submanifolds with codimension one the external rotation tensor vanishes by definition. The curvature associated to the external rotation tensor is known as the outer curvature $\Omega^{I}{}_{Jab}$ and is given by
$\Omega^{I}_{Jab}=2\partial_{[a}\omega_{b]}{}^{I}{}_{J}-2\omega_{[a|}{}^{I}{}_{K}\omega_{|b]}{}^{K}{}_{J}.$ (54)
The set of submanifold curvatures $\{\mathcal{R}^{a}{}_{bcd},K_{ab}{}^{I},\Omega^{I}_{Jab}\}$ are related to the ambient space curvature $R^{\rho}{}_{\sigma\mu\nu}$ through the integrability conditions known as the Codazzi-Mainardi equation
$\mathcal{D}_{a}\tilde{K}_{bc}{}^{I}-\mathcal{D}_{b}\tilde{K}_{ac}{}^{I}=R^{I}{}_{cab}+\hat{\upsilon}^{d}F_{ab}\tilde{K}_{dc}{}^{I},$ (55)
the Gauss-Codazzi equation
$\mathcal{R}^{d}_{cab}=R^{d}_{cab}+\left(\tilde{K}_{bc}{}^{I}\tilde{K}_{af}{}^{J}-\tilde{K}_{ac}{}^{I}\tilde{K}_{bf}{}^{J}\right)h^{fd}\delta_{IJ},$ (56)
and the Ricci-Voss equation
$\Omega^{I}{}_{Jab}=-R^{I}_{Jab}-2h^{cd}\tilde{K}_{[a|c}{}^{I}\tilde{K}_{|b]d}{}^{J}\delta_{IJ},$ (57)
where in (55)-(57) we defined
$\tilde{K}_{ab}{}^{I}\equiv K_{ab}{}^{I}-\frac{1}{2}\hat{\upsilon}^{I}F_{ab}=n^{I}_{\mu}D_{a}u^{\mu}_{b}.$ (58)
III Currents and conservation equations
Let’s consider the action of a (d-1)-dimensional submanifold $\Sigma$ of a d-dimensional manifold $\mathcal{M}_{d}$
$S=S_{\Sigma}\left[\hat{\phi},\bar{h}_{ab},\tau_{a},K_{ab}{}^{I}\right],$ (59)
where $\phi$ are the fields describing the microscopic theory with $\hat{\phi}$ their value at the sub-manifold. A variation of this action will take the form [^6] Notice that there is no transverse index on $\mathcal{D}$ or $\delta K$ as the boundary is a codimension 1 surface.
$\delta S$ $=\int_{\partial\mathcal{M}}d^{d-1}\sigma\hat{e}\left(\left(\mathcal{T}_{S}\right)^{a}\delta\tau_{a}+\frac{1}{2}\left(\mathcal{T}_{S}\right)^{ab}\delta\bar{h}_{ab}+\mathcal{D}^{ab}\delta K_{ab}\right),$ (60)
where $\hat{e}=\sqrt{\frac{\det\bar{h}_{ab}}{2\hat{\Phi}}}$, $\mathcal{T}^{a}_{S}$ is the surface boundary current, $\mathcal{T}^{ab}_{S}$ is the surface Cauchy stress-mass tensorGeracie:2016 , and $\mathcal{D}^{ab}$ is the bending moment encoding elastic responses. The equations of motion can be derived by performing a Lagrangian variation of the actionArmas:2020 , such variation transforms the bulk fields and keeps the embedding fixed. Namely, the Lagrangian variations under diffeomorphisms and U(1) mass transformations take the explicit form
$\delta\tau_{a}$ $=\tau_{\rho}D_{a}\xi^{\rho}-F_{a\rho}\xi^{\rho}\,,$ (61)
$\delta\bar{h}_{ab}$ $=2\bar{h}_{\rho(a}D_{b)}\xi^{\rho}-2\xi^{\sigma}\left[\bar{\mathcal{K}}_{\sigma(a}\tau_{b)}+\tau_{\sigma}\tau_{(a}D_{b)}\Phi+\tau_{a}\tau_{b}\nabla_{\sigma}\Phi+F_{\sigma(a}\tau_{b)}\Phi\right]-2\tau_{(a}\partial_{b)}\sigma\,,$ (62)
$\delta K_{ab}$ $=n_{\alpha}D_{a}D_{b}\xi^{\alpha}-R^{\alpha}{}_{ba\lambda}\xi^{\lambda}n_{\alpha}+\left(F_{\lambda b}n_{\beta}D_{a}\hat{\upsilon}^{\beta}+vu^{\beta}_{b}D_{a}F_{\beta\lambda}+vF_{\mu\lambda}n^{\mu}K_{ab}\right)\xi^{\lambda}$ (63)
$\hphantom{=}+vF_{b\alpha}D_{a}\xi^{\alpha}-\frac{1}{2}vF_{ab}n^{\alpha}n_{\lambda}D_{\alpha}\xi^{\lambda},$
where $\bar{\mathcal{K}}_{\mu\nu}=-\frac{1}{2}\mathcal{L}_{\hat{\upsilon}}\bar{h}_{\mu\nu}$. Invariance of (60) under this set of transformations gives out the derivative constraint
$F_{ab}=0\,,$ (64)
the tangential equations of motion
$D_{a}\left[\left(\mathcal{T}_{S}\right)^{ab}\tau_{a}\right]$ $=0\,,$ (65)
$D_{a}\left(\mathcal{T}_{S}\right)^{a}-\mathcal{K}^{\Sigma}_{ab}\left(\mathcal{T}_{S}\right)^{ab}-2\left(\mathcal{T}_{S}\right)^{ab}\tau_{b}D_{a}\hat{\Phi}$ $=2\hat{\upsilon}^{d}D_{a}\left(\mathcal{D}^{ab}K_{db}\right)-\hat{\upsilon}^{a}\mathcal{D}^{cd}D_{a}K_{cd},$ (66)
$D_{a}\left(\mathcal{T}_{S}\right)^{ac}$ $=h^{ca}\mathcal{D}^{cd}D_{a}K_{cd}-2h^{cd}D_{a}\left(\mathcal{D}^{ab}K_{db}\right),$ (67)
and the transverse constraint
$K_{ab}\left(\mathcal{T}_{S}\right)^{ab}$ $=D_{a}D_{b}\mathcal{D}^{ab}-h^{cd}\mathcal{D}^{ab}K_{ac}K_{bd}+\mathcal{D}^{ab}R_{ab}$ (68)
$+f_{a}\left[\left(\mathcal{T}_{S}\right)^{a}+vD_{b}\mathcal{D}^{ab}+\hat{\upsilon}^{b}K_{bd}\mathcal{D}^{ad}+2\mathcal{D}^{ab}D_{b}v\right],$
where we defined $\mathcal{K}^{\Sigma}_{ab}=-\frac{1}{2}\mathcal{L}_{\hat{\upsilon}}\bar{h}_{ab},R_{ab}=R^{\lambda}{}_{a\mu b}n^{\mu}n_{\lambda},$ and $f_{a}=F_{a\lambda}n^{\lambda}$. This last equation is the shape equation and will determine the embedding function $n_{\mu}X^{\mu}$.
IV Equilibrium Constituve Relations
Our focus will be on writing out the membrane constitutive relations on equilibrium configurations. Such configurations are characterized by the existence of a time-like killing vector $V^{\mu}$ that characterizes the time independence of the fluid configuration. In particular to have an equilibrium configuration the background should be invariant under the symmetry transformation
$\mathcal{L}_{V}\tau_{\mu}=0\qquad\text{and}\qquad\mathcal{L}_{V}\bar{h}_{\mu\nu}=2\tau_{(\mu}\partial_{\nu)}\Lambda^{V},$ (69)
where $\Lambda^{V}$ is the associated gauge parameter. The equivalent statement should hold for the pullback onto the surface $k^{a}=u^{a}_{\mu}k^{\mu}$, namely
$\mathcal{L}_{V}\tau_{a}=0\qquad\text{and}\qquad\mathcal{L}_{V}\bar{h}_{ab}=2\tau_{(a}\partial_{b)}\Lambda^{K}.$ (70)
An equilibrium state is specified by a given set of gauge invariant physical parameters $\{T,\mu,u^{\rho}\}$ identified with the temperature, mass chemical potential, and four velocity respectively. In equilibrium the associated generating function $\mathcal{F}[T,\mu,u]$ should be a function of the background geometry and the symmetry parameter set $\{V^{\mu},\Lambda^{V}\}$. This means the physical parameters have a solution in terms of the background geometry and symmetry set, such solution can easily be found to be Armas:2020 ; Jain:2020
$u^{a}=\frac{V^{a}}{V^{c}\tau_{c}}\,,\qquad T=\frac{T_{0}}{V^{c}\tau_{c}}\,,\qquad\frac{\mu}{T}=-\frac{1}{2}\frac{\Lambda^{V}}{T_{0}}+\frac{1}{2T}\bar{h}_{cd}u^{c}u^{d},$ (71)
where $T_{0}$ is a reference temperature, and $u^{\mu}$ has been normalised to $u^{\mu}\tau_{\mu}=1$. All we have to do now is write down the most general generating function $\mathcal{F}[T,\mu,K]$ up to an arbitrary number of derivatives. The equilibrium conditions (70) give rise to the conditions
$D_{a}T$ $=0\,,$ (72)
$u^{a}D_{a}\mu$ $=0\,,$ (73)
$h^{rs}D_{s}\mu+a^{r}$ $=0\,,$ (74)
$\theta$ $=0\,,$ (75)
$\sigma^{ab}$ $=0\,,$ (76)
where we defined $\Delta^{a}_{c}=h^{ad}\bar{h}_{dc}$ and where we used the identifications
$a^{d}$ $=u^{c}D_{c}u^{d}\,,$ (77)
$\theta$ $=\Delta^{b}_{c}D_{b}u^{c}\,,$ (78)
$\sigma^{fg}$ $=\left[h^{bg}\Delta^{f}_{c}+h^{bf}\Delta^{g}_{d}-h^{fg}\frac{\Delta^{b}_{c}}{d-2}\right]D_{b}u^{c}\,,$ (79)
$\Omega^{fg}$ $=\frac{1}{2}\left(h^{bg}\Delta^{f}_{c}-h^{bf}\Delta^{g}_{c}\right)D_{b}u^{c}\,,$ (80)
with $a^{d}$ as the acceleration, $\theta$ as the compresibility, $\sigma^{fg}$ as the shear, and $\Omega^{fg}$ as the vorticity. To finalize we can note the following definitions that will come in handy later
$\bar{u}_{a}$ $\equiv\bar{h}_{ab}u^{b}\,,$ (81)
$\bar{a}_{c}$ $\equiv\bar{h}_{cd}a^{d}\,,$ (82)
$\bar{u}^{2}$ $\equiv\bar{h}_{ab}u^{a}u^{b}=\bar{u}_{a}u^{a}\,,$ (83)
$\bar{a}^{2}$ $\equiv\bar{h}_{cd}a^{c}a^{d}.$ (84)
IV.1 Equilibrium Partition Function
We need to classify all possible gauge invariant scalars that can be constructed using $\{\mu,T,u^{a},K_{ab},\bar{h}_{cd},\tau_{d}\}$ up to second order in derivatives [^7] The Canham-Helfrich functional used to derive the biconcave shape of blood cells contains terms of quadratic order in the extrinsic curvature meaning that for consistency we must keep all terms up to the same order in derivatives.. We are assuming both $\{T,\mu\}$ are zeroth order while $K_{ab}$ is taken to be first order.
Order zero hydrostatic scalars
$T$ $\mu$
Order one hydrostatic scalars
$F_{(1)}=K_{ab}h^{ab}$ $F_{(2)}=K_{ab}u^{a}u^{b}$
Order two hydrostatic scalars
$S_{(1)}=\left(K_{ab}h^{ab}\right)\left(K_{cd}h^{cd}\right)$ $S_{(2)}=\left(K_{ab}K_{cd}h^{ac}h^{bd}\right)$ $S_{(3)}=\left(K_{ab}K_{cd}u^{a}h^{bc}u^{d}\right)$ $S_{(4)}=\left(K_{ab}u^{a}u^{b}\right)\left(K_{cd}u^{c}u^{d}\right)$
$S_{(5)}=a^{c}a^{d}\bar{h}_{cd}$ $S_{(6)}=\left(K_{cd}u^{d}\right)a^{c}$ $S_{(7)}=\Omega^{ab}\Omega^{cd}\bar{h}_{ac}\bar{h}_{bd}$
Table 1: Summary of zeroth, first and second order independent inequivalent scalars which may contribute to the equilibrium partition function.
Where we assumed zero torsion and a flat ambient spacetime, ignore terms related to each other through partial integration, and used the corresponding integrability condition
$\mathcal{R}^{d}_{cab}$ $=\left(K_{bc}K_{af}-K_{ac}K_{bf}\right)h^{fd}\,,$ (85)
to ignore any explicit curvature terms. The equilibrium partition function can then be written as
$\mathcal{F}$ $=\int d^{d-1}\sigma\hat{e}\left[\chi(T,\mu)+\sum_{i=1}^{2}a^{(1)}_{i}(T,\mu)F_{(i)}+\sum_{i=1}^{7}a^{(2)}_{i}(T,\mu)S_{(i)}\right].$ (86)
Where $\chi(T,\mu)$ is the surface tension, and $\{a^{(1)}_{i}(T,\mu),a^{(2)}_{i}(T,\mu)\}$ are generic functions of the temperature and chemical potential. From this point onward, and for the sake of simplicity, we will omit the functional dependence of these aforementioned functions. The variation of the equilibrium partition function will take the form
$\delta\mathcal{F}$ $=\int d^{d-1}\sigma\hat{e}\left(\frac{h^{ab}\delta\bar{h}_{ab}-2\hat{\upsilon}^{a}\delta\tau_{a}}{2}\right)\left[\chi+\sum_{i=1}^{2}a^{(1)}_{i}F_{(i)}+\sum_{i=1}^{7}a_{i}^{(2)}S_{(i)}\right]$ (87)
$+\int d^{d-1}\sigma\hat{e}\left[s\delta T+n\delta\mu+\sum_{i=1}^{2}\left(s^{(1)}_{i}\delta T+n^{(1)}_{i}\delta\mu\right)F_{(i)}+\sum_{i=1}^{7}\left(s_{i}^{(2)}\delta T+n_{i}^{(2)}\delta\mu\right)S_{(i)}\right]$
$+\int d^{d-1}\sigma\hat{e}\left[\sum_{i=1}^{2}a^{(1)}_{i}\delta F_{(i)}+\sum_{i=1}^{7}a^{(2)}_{i}\delta S_{(i)}\right],$
where we define surface entropy density $s$, surface particle density $n$, and the surface susceptibilities $\{s^{(I)}_{i},n^{(I)}_{i}\}$
$s$ $=\left(\frac{\partial\chi}{\partial T}\right)_{\mu},\qquad$ $s^{(I)}_{i}$ $=\left(\frac{\partial a^{(I)}_{i}}{\partial T}\right)_{\mu},$ (88)
$n$ $=\left(\frac{\partial\chi}{\partial\mu}\right)_{T},\qquad$ $n^{(I)}_{i}$ $=\left(\frac{\partial a^{(I)}_{i}}{\partial\mu}\right)_{T},$
with $I=\{1,2\}$, and where we used the identity
$\delta\hat{e}$ $=\frac{\hat{e}}{2}\left(h^{ab}\delta\bar{h}_{ab}-2\upsilon^{a}\delta\tau_{a}\right).$ (89)
To find the relevant currents we will use the following variations of (71), (77), and (80)
$\delta T$ $=-Tu^{a}\delta\tau_{a}\,,$ (90)
$\delta\mu$ $=-\left(\mu+\frac{1}{2}\bar{u}^{2}\right)u^{a}\delta\tau_{a}+\frac{1}{2}u^{a}u^{b}\delta h_{ab}\,,$ (91)
$\delta u^{a}$ $=-u^{a}u^{b}\delta\tau_{b}\,,$ (92)
$\delta a^{d}$ $=\delta u^{c}D_{c}u^{d}+u^{c}\left(D_{c}\delta u^{d}+\delta\Gamma^{d}_{cf}u^{f}\right)=-\delta h^{cd}D_{c}\mu-h^{cd}D_{c}\delta\mu\,,$ (93)
$\delta\Omega^{fg}\Omega^{rs}\bar{h}_{fr}\bar{h}_{gs}$ $=\frac{1}{2}\left[\bar{u}_{a}\bar{h}_{bc}\Omega^{ab}-D_{c}\left(\mu+\frac{1}{2}\bar{u}^{2}\right)\right]\Omega^{cd}\delta\tau_{d}+\frac{1}{2}\left[\Omega^{af}\Omega^{bg}\bar{h}_{ab}\bar{h}_{fg}\right]u^{a}\delta\tau_{a}$ (94)
$-\frac{1}{2}\delta\bar{h}_{cd}\Omega^{cf}\Omega^{dg}\bar{h}_{fg}-\frac{1}{2}\bar{h}_{bc}\Omega^{cf}u^{b}u^{d}D_{f}\delta\tau_{d}-\frac{1}{2}\Omega^{cd}u^{b}D_{c}\delta h_{bd}\,,$
$\delta\Gamma^{a}_{cd}$ $=-\hat{\upsilon}^{a}D_{c}\delta\tau_{d}+\frac{1}{2}\delta h^{ab}\left(2D_{(c}\bar{h}_{d)b}-D_{b}\bar{h}_{cd}\right)+\frac{1}{2}h^{ab}\left(D_{(c}\delta\bar{h}_{d)b}-D_{b}\delta\bar{h}_{cd}\right),$ (95)
as well as the geometric identities
$\delta h^{ab}$ $=-h^{ac}h^{bd}\delta\bar{h}_{cd}+2\hat{\upsilon}^{(a}h^{b)c}\delta\tau_{c}\,,$ (96)
$\delta\upsilon^{a}$ $=\left(2\check{\Phi}h^{ab}+\hat{\upsilon}^{a}\hat{\upsilon}^{b}\right)\delta\tau_{b}-h^{ac}\hat{\upsilon}^{d}\delta\bar{h}_{cd}.$ (97)
Using (60) the surface energy current is found to be
$\left(\mathcal{T}_{\mathcal{S}}\right)^{c}$ $=-\chi\hat{\upsilon}^{c}-\left(\epsilon+\chi+\frac{n}{2}\bar{u}^{2}\right)u^{c}-\left(\sum_{i=1}^{2}\zeta^{(1)c}F_{(i)}+\sum_{i=1}^{7}\zeta^{(1)c}S_{(i)}\right)$ (98)
$+\left[2\left(a_{1}^{(1)}+2a_{1}^{(2)}F_{(1)}\right)\hat{\upsilon}^{a}K_{ab}+4a_{2}^{(2)}\hat{\upsilon}^{a}\left(K_{ar}h^{rs}K_{sb}\right)+2a_{3}^{(2)}\hat{\upsilon}^{a}\left(K_{ar}u^{r}\right)\left(K_{db}u^{d}\right)\right]h^{bc}$
$-\left[4a^{(2)}_{5}\hat{a}_{f}D_{g}\mu+2a^{(2)}_{6}K_{af}u^{a}D_{g}\mu\right]\upsilon^{(f}h^{g)c}$
$-\left[2a^{(2)}_{5}h^{fg}D_{f}D_{g}\mu+h^{bf}D_{f}\left(a^{(2)}_{6}K_{ab}u^{a}\right)\right]\left(\mu+\frac{1}{2}\bar{u}^{2}\right)u^{c}$
$+\left[2a^{(2)}_{(7)}\bar{u}_{a}\bar{h}_{bd}\Omega^{ab}-2a^{(2)}_{(7)}D_{d}\left(\mu+\frac{1}{2}\bar{u}^{2}\right)\right]\Omega^{cd}$
$+\left[2\left(a_{2}^{(1)}+2a_{4}^{(2)}F_{(2)}\right)K_{ab}u^{a}u^{b}-2a_{3}^{(2)}\left(K_{ar}h^{rs}K_{sb}\right)u^{a}u^{b}+a^{(2)}_{(6)}K_{fg}a^{f}u^{g}\right]u^{c}$
$+\left[D_{f}\left(a^{(2)}_{(7)}\bar{h}_{bc}\Omega^{cf}u^{b}\right)-a^{(2)}_{(7)}\Omega^{af}\Omega^{bg}\bar{h}_{ab}\bar{h}_{fg}\right]u^{c},$
the surface Cauchy stress-mass tensor is found to be
$\left(\mathcal{T}_{\mathcal{S}}\right)^{cd}$ $=\chi h^{cd}+nu^{c}u^{d}+\left(\sum_{i=1}^{2}\xi^{(1)cd}_{i}F_{(i)}+\sum_{i=1}^{7}\xi^{(2)cd}_{i}S_{(i)}\right)$ (99)
$-\left[2\left(a_{1}^{(1)}+2a_{1}^{(2)}F_{(1)}\right)K_{ab}+4a_{2}^{(2)}\left(K_{ar}h^{rs}K_{sb}\right)+2a_{3}^{(2)}\left(K_{ar}u^{r}\right)\left(K_{bs}u^{s}\right)\right]h^{ac}h^{bd}$
$+\left[4a^{(2)}_{5}a_{f}D_{g}\mu+2a^{(2)}_{6}K_{af}u^{a}D_{g}\mu\right]h^{cf}h^{dg}+\left[a^{(2)}_{5}2h^{fg}D_{f}a_{g}+a^{(2)}_{6}h^{bf}D_{f}\left(K_{ab}u^{a}\right)\right]u^{c}u^{d}$
$+\left[2a^{(2)}_{7}\Omega^{ra}\Omega^{sb}\bar{h}_{rf}\bar{h}_{sg}\bar{h}_{ab}\right]h^{cf}h^{dg}-2D_{b}\left(a^{(2)}_{7}\Omega^{b(c}u^{d)}\right),$
and the bending moment is given by
$\mathcal{D}^{cd}$ $=a_{1}^{(1)}h^{cd}+a_{2}^{(1)}u^{c}u^{d}+\mathcal{Y}^{cdab}K_{ab}+a^{(2)}_{6}a^{c}u^{d},$ (100)
with $\mathcal{Y}^{cdab}$ the Young modulus
$\mathcal{Y}^{cdab}$ $=2a_{1}^{(2)}h^{cd}h^{ab}+2a_{2}^{(2)}h^{ac}h^{bd}+2a_{3}^{(2)}u^{(c}h^{d)b}u^{a}+2a_{4}^{(2)}u^{c}u^{d}u^{a}u^{b}.$ (101)
Where we have defined the internal energy $\epsilon$ as Armas:2020
$\epsilon=-\chi+Ts+n\mu,$ (102)
and the tensors $\{\zeta^{(I)c}_{i},\xi^{(I)cd}_{i}\}$ as
$\zeta^{(a)c}_{i}$ $=a^{(I)}_{i}\hat{\upsilon}^{c}+L^{(I)}_{i}u^{c}\,,$ (103)
$\xi^{(a)cd}_{i}$ $=a^{(I)}_{i}h^{cd}+n^{(I)}_{i}u^{c}u^{d}\,,$ (104)
with
$L^{(I)}_{i}$ $=s_{i}^{(I)}T+n_{i}^{(I)}\mu+\frac{1}{2}n_{i}^{(I)}\bar{u}^{2}.$ (105)
Using (88), the relation (102) is equivalent to the thermodynamics first law
$d\epsilon$ $=Tds+\mu dn$ (106)
IV.2 Axisymmetric configurations in 3+1 flat ambient spacetime
For applications concerning biological systems, such as the lipid bilayer membranes discussed in Helfrich ; Armas:2020 . It is enough to restrict the ambient spacetime geometry to
$\tau_{\mu}=\delta_{\mu}^{t}\,,\qquad h_{\mu\nu}$ $=\delta^{i}_{\mu}\delta_{\nu}^{j}\delta_{ij}\,,\qquad\hat{\upsilon}^{\rho}=-\delta^{\rho}_{t}\,,\qquad h^{\mu\nu}=\delta^{\mu}_{i}\delta^{\nu}_{j}\delta^{ij}\,,\quad m_{\mu}=0,$ (107)
where we are assuming a time direction $t$ and a Cartesian three-dimensional space with coordinates $x^{i}=(x,y,z)$. The equilibrium condition on this background, with gauge parameter $\Lambda^{K}=0$, has a solution for the Killing vector $V^{\mu}$ given by
$V=c_{0}\partial_{t}+c_{1}\left(z\partial_{y}-y\partial_{z}\right)+c_{2}\left(x\partial_{z}-z\partial_{x}\right)+c_{3}\left(y\partial_{x}-x\partial_{y}\right),$ (108)
where $\{c_{0},c_{1},c_{2},c_{3}\}$ are constants that can be chosen at our convenience. We will be interested in studying membranes subjected to axysimmetric flows, therefore we need to set $c_{1}=c_{2}=0$. This corresponds to a flow on the ambient spacetime with velocity
$u^{\mu}$ $=\delta^{\mu}_{t}+\frac{\omega}{2}\left(x\delta^{\mu}_{y}-y\delta^{\mu}_{x}\right),$ (109)
where we have defined $\omega=-\frac{2c_{3}}{c_{0}}$. Notice that this flow corresponds to a flow with a constant spatial vorticity vector $\boldsymbol{\omega}=(0,0,\omega)$. Membranes with axisymmetric symmetry can be parametrized through the embedding mapArmas:2020 :
[…]
The coordinates on the membrane will then be $(t,\rho,\phi)$. Under this embedding it is easy to find that
$\tau_{a}=\delta^{t}_{a}\,,\quad\bar{h}_{ab}$ $=\left(\sec^{2}\psi\right)\delta^{\rho}_{a}\delta^{\rho}_{b}+\rho^{2}\delta^{\phi}_{a}\delta^{\phi}_{b}\,,\quad\hat{\upsilon}^{a}=-\delta^{a}_{t}\,,\quad h^{ab}=\left(\cos^{2}\psi\right)\delta^{a}_{\rho}\delta^{b}_{\rho}+\frac{\delta^{a}_{\phi}\delta^{b}_{\phi}}{\rho^{2}}.$ (111)
The associated fluid velocity $u^{a}$ will be given by
$u^{a}=\delta^{a}_{t}+\frac{\omega}{2}\delta^{a}_{\phi}.$ (112)
The acceleration, compressibility, shear, and vorticity will be given by
$a^{b}$ $=\left(-\frac{\omega^{2}}{4}\rho\cos^{2}\psi\right)\delta^{b}_{\rho}\,,$ (113)
$\theta$ $=0\,,$ (114)
$\sigma^{ab}$ $=0\,,$ (115)
$\Omega^{ab}$ $=\left(\frac{\omega}{2\rho}\cos^{2}\psi\right)\left(\delta^{a}_{\rho}\delta^{b}_{\phi}-\delta^{a}_{\phi}\delta^{b}_{\rho}\right).$ (116)
Notice that the equilibrium condition implies that temperature is constant and related to $c_{0}$ through $\frac{T_{0}}{c_{0}}$ while the chemical potential $\mu$ will be a function of $\rho$ given by
$\mu(\rho)=\frac{\omega^{2}\rho^{2}}{8}.$ (117)
The extrinsic curvature on this configuration is given by
$K_{ab}=-\left(\psi^{\prime}\sec\psi\right)\delta^{\rho}_{a}\delta^{\rho}_{b}-\left(\sin\psi\right)\delta_{a}^{\phi}\delta_{b}^{\phi}.$ (118)
Using all this information, we show in table 2 all the scalars from table 1 in terms of the embedding variables. Before finishing this section, notice that for any function $f(T,\mu)$, it will be true under the equilibrium condition that
$D_{b}f$ $=\left(\frac{\partial f}{\partial\mu}\right)D_{b}\mu,$ (119)
$D_{a}D_{b}f$ $=\left(\frac{\partial^{2}f}{\partial\mu^{2}}\right)\left(D_{a}\mu\right)\left(D_{b}\mu\right)+\left(\frac{\partial f}{\partial\mu}\right)D_{a}D_{b}\mu.$ (120)
Order zero hydrostatic scalars
$T$
$\mu=\frac{\omega^{2}\rho^{2}}{8}$
Order one hydrostatic scalars
$F_{(1)}=K_{ab}h^{ab}=-\frac{\sin\psi}{\rho}-\psi^{\prime}\cos\psi$
$F_{(2)}=K_{ab}u^{a}u^{b}=-\frac{1}{4}\omega^{2}\rho\sin\psi$
Order two hydrostatic scalars
$S_{(1)}=\left(K_{ab}h^{ab}\right)\left(K_{cd}h^{cd}\right)=\frac{\left(\psi^{\prime}\cos\psi+\sin\psi\right)^{2}}{\rho^{2}}$
$S_{(2)}=\left(K_{ab}K_{cd}h^{ac}h^{bd}\right)=\frac{\sin^{2}\psi}{\rho^{2}}+\left(\psi^{\prime}\right)^{2}\cos^{2}\psi$
$S_{(3)}=\left(K_{ab}K_{cd}u^{a}h^{bc}u^{d}\right)=\frac{1}{4}\omega^{2}\sin^{2}\psi$
$S_{(4)}=\left(K_{ab}u^{a}u^{b}\right)\left(K_{cd}u^{c}u^{d}\right)=\frac{1}{16}\omega^{4}\rho^{2}\sin^{2}\psi$
$S_{(5)}=a^{c}a^{d}\bar{h}_{cd}=\frac{1}{16}\omega^{4}\rho^{2}\cos^{2}\psi$
$S_{(6)}=\left(K_{cd}u^{d}\right)a^{c}=0$
$S_{(7)}=\Omega^{ab}\Omega^{cd}\bar{h}_{ac}\bar{h}_{bd}=\frac{1}{2}\omega^{2}\cos^{2}\psi$
Table 2: Summary of zeroth, first and second order independent inequivalent scalars for an axisymmetric system.
V Young-Laplace equation for axisymmetric solutions
The equation describing the shape of the membrane is given by (68) and can be considered a generalization of Young-Laplace equation. In flat ambient spacetime, it takes the form
$D_{a}D_{b}\mathcal{D}^{ab}-\left(\mathcal{T_{S}}\right)^{ab}K_{ab}-h^{cd}\mathcal{D}^{ab}K_{ac}K_{bd}=0.$ (121)
In this section, we will focus on the immediate generalization of the case studied in Armas:2020 . This generalization will consist of keeping only the terms in the equilibrium partition function that depend only on the vorticity through their chemical potential dependence. This means keeping $\chi,a^{(1)}_{1},a^{(2)}_{1},$ and $a^{(2)}_{2}$ as arbitrary functions of the temperature and chemical potential while setting all other $a^{I}_{i}$ coefficients to zero [^8] The explicit form of the full Young-Laplace equation taking all coefficients into account is quite cumbersome and by itself it is not very instructive, so we will leave its full analysis for future work.. Therefore, the equilibrium partition function will simply look like
$\mathcal{F}=\int d^{3}\sigma e\left[\chi+a_{1}^{(1)}K+a_{1}^{(2)}K^{2}+a_{2}^{(2)}\left(K\cdot K\right)\right],$ (122)
where we defined the traces
$K_{1}$ $=K_{ab}h^{ba}\,,$ (123)
$K_{2}$ $=K_{ab}h^{bc}K_{cd}h^{da}.$ (124)
The relevant constitutive relations will take the form
$\left(\mathcal{T_{S}}\right)^{cd}$ $=\left[\chi+a_{1}^{(1)}K_{1}+a_{1}^{(2)}K_{1}^{2}+a_{2}^{(2)}K_{2}\right]h^{cd}$ (125)
$+\left[n+n^{(1)}_{1}K_{1}+n^{(2)}_{1}K_{1}^{2}+n^{(2)}_{2}K_{2}\right]u^{c}u^{d}$
$-2\left(a^{(1)}_{1}+2a^{(2)}_{1}K_{1}\right)h^{ac}h^{bd}K_{ab}-4a^{(2)}_{2}h^{rs}h^{ac}h^{bd}K_{ar}K_{sb},$
$\mathcal{D}^{cd}$ $=a_{1}^{(1)}h^{cd}+2a_{1}^{(2)}h^{cd}K_{1}+2a_{2}^{(2)}h^{ac}h^{bd}K_{ab}.$ (126)
We can first compute the parts that appear in the Young-Laplace equation without covariant derivatives:
$\left(\mathcal{T}_{S}\right)^{ab}K_{ab}$ $=\chi K_{1}+a^{(1)}_{1}K_{1}^{2}+a_{1}^{(2)}K_{1}^{3}+a_{2}^{(2)}K_{1}K_{2}+nk_{0}+n_{1}^{(1)}K_{1}k_{0}$ (127)
$+n_{1}^{(2)}K_{1}^{2}k_{0}+n^{(2)}_{2}K_{2}k_{0}-2a_{1}^{(1)}K_{2}-4a_{1}^{(2)}K_{1}K_{2}-4a_{2}^{(2)}K_{3},$
$h^{cd}\mathcal{D}^{ab}K_{ac}K_{bd}$ $=a_{1}^{(1)}K_{2}+2a_{1}^{(2)}K_{1}K_{2}+2a_{2}^{(2)}K_{3},$ (128)
where
$k_{c}$ $=K_{cd}u^{d}\,,$ (129)
$k_{0}$ $=k_{c}u^{c}\ ,$ (130)
$K_{3}$ $=K_{ab}h^{bc}K_{cd}h^{dr}K_{rs}h^{sa}.$ (131)
Combining these two terms, we end up with
$\left(\mathcal{T_{S}}\right)^{cd}+h^{cd}\mathcal{D}^{ab}K_{ac}K_{bd}$ $=\chi K_{1}+a_{1}^{(1)}K_{1}^{2}-a_{1}^{(1)}K_{2}+\left(a_{1}^{(2)}+a_{2}^{(2)}\right)\left(K_{1}^{2}-2K_{2}\right)K_{1}$ (132)
$+nk_{0}+n_{1}^{(1)}K_{1}k_{0}+n_{1}^{(2)}K^{2}_{1}k_{0}+n_{2}^{(2)}K_{2}k_{0},$
where we used the following identity between traces
$K_{3}=\frac{3}{2}K_{1}K_{2}-\frac{1}{2}K_{1}^{3}.$ (133)
We can now take a look at the term with derivatives of the bending moment
$D_{c}D_{d}\mathcal{D}^{cd}$ $=h^{cd}D_{c}D_{d}a_{1}^{(1)}+2h^{cd}\left[K_{1}D_{c}D_{d}a_{1}^{(2)}+2\left(D_{c}K_{1}\right)\left(D_{d}a_{1}^{(2)}\right)+a_{1}^{(2)}D_{c}D_{d}K_{1}\right]$ (134)
$+2h^{ac}h^{bd}\left[K_{ab}D_{c}D_{d}a^{(2)}_{2}+2\left(D_{c}K_{ab}\right)\left(D_{d}a^{(2)}_{2}\right)+a^{(2)}_{2}D_{c}D_{d}K_{ab}\right]$
$=2\left(a^{(2)}_{1}+a_{2}^{(2)}\right)D^{2}K_{1}+\left(\beta^{(1)}_{1}+2\beta^{(2)}_{1}K_{1}\right)\bar{a}^{2}-\left(n^{(1)}_{1}+2n^{(2)}_{1}K_{1}\right)\left(D\cdot a\right)$
$+2\beta^{(2)}_{2}\left(K_{cd}a^{c}a^{d}\right)-2n^{(2)}_{2}\left(K_{cd}h^{ac}D_{a}a^{d}\right)-4\left(n^{(2)}_{1}+n^{(2)}_{2}\right)a^{c}D_{c}K_{1},$
where we used the Codazzi-Mainardi equation and where we defined
$\beta^{I}_{(i)}$ $=\frac{\partial n}{\partial\mu}\,,$ (135)
$D^{2}$ $=h^{cd}D_{c}D_{d}\,,$ (136)
$D\cdot a$ $=D_{c}a^{c}.$ (137)
The Young-Laplace equation takes the form
$2\left(a^{(2)}_{1}+a_{2}^{(2)}\right)D^{2}K_{1}+\left(\beta^{(1)}_{1}+2\beta^{(2)}_{1}K_{1}\right)\bar{a}^{2}-\left(n^{(1)}_{1}+2n^{(2)}_{1}K_{1}\right)\left(D\cdot a\right)$ (138)
$+2\beta^{(2)}_{2}\left(K_{cd}a^{c}a^{d}\right)-2n^{(2)}_{2}\left(K_{cd}h^{ac}D_{a}a^{d}\right)-4\left(n^{(2)}_{1}+n^{(2)}_{2}\right)a^{c}D_{c}K_{1}$
$-\chi K_{1}-a_{1}^{(1)}K_{1}^{2}+a_{1}^{(1)}K_{2}-\left(a_{1}^{(2)}+a_{2}^{(2)}\right)\left(K_{1}^{2}-2K_{2}\right)K_{1}$
$+nk_{0}+n_{1}^{(1)}K_{1}k_{0}+n_{1}^{(2)}K^{2}_{1}k_{0}+n_{2}^{(2)}K_{2}k_{0}$ $=0.$
From the form of the Young-Laplace equation we can notice that it is convenient to change the transport coefficients from the set $\{\chi,a^{(1)}_{1},a^{(2)}_{1},a^{(2)}_{2}\}$ to the set $\{\chi,a^{(1)}_{1},a^{(2)}_{2},\kappa\}$, where $\kappa$ is defined through
$a^{(2)}_{1}$ $=\kappa-a_{2}^{(2)}.$ (139)
The corresponding derivatives of $\kappa$ with respect to $\mu$ are related to $n^{(2)}_{1},n^{(2)}_{2},\beta^{(2)}_{1}$ and $\beta^{(2)}_{2}$ through
$n^{(2)}_{1}$ $=n_{\kappa}-n_{2}^{(2)}\,,$ (140)
$\beta^{(2)}_{1}$ $=\beta_{\kappa}-\beta_{2}^{(2)}.$ (141)
Under this redefinition, the Young-Laplace equation becomes the following
$2\kappa D^{2}K_{1}+\left(\beta^{(1)}_{1}+2\left(\beta_{\kappa}-\beta^{(2)}_{2}\right)K_{1}\right)\bar{a}^{2}-\left(n^{(1)}_{1}+2\left(n_{\kappa}-n_{2}^{(2)}\right)K_{1}\right)\left(D\cdot a\right)$ (142)
$+2\beta^{(2)}_{2}\left(K_{cd}a^{c}a^{d}\right)-2n^{(2)}_{2}\left(K_{cd}h^{ac}D_{a}a^{d}\right)-4n_{\kappa}a^{c}D_{c}K_{1}$
$-\chi K_{1}a_{1}^{(1)}\left(K_{2}-K_{1}^{2}\right)-\kappa\left(K_{1}^{2}-2K_{2}\right)K_{1}$
$+nk_{0}+n_{1}^{(1)}K_{1}k_{0}+n_{2}^{(2)}\left(K_{2}-K^{2}_{1}\right)k_{0}+n_{\kappa}K_{1}^{2}k_{0}$ $=0.$
For simplicity we will restate the Young-Laplace equation as
$Y+A_{\beta}+N^{(1)}_{1}+N^{(2)}_{2}+N_{\kappa}-\frac{n}{4}\omega^{2}\rho\sin\psi=0,$ (143)
where $Y$ contain the terms that do not involve the vorticity explicitly
$Y$ $\equiv 2\kappa D^{2}K_{1}-\chi K_{1}+a_{1}^{(1)}\left(K_{2}-K^{2}_{1}\right)-\kappa\left(K^{2}_{1}-2K_{2}\right)K_{1}$ (144)
$=-\frac{2a_{1}^{(1)}\psi^{\prime}\cos\psi\sin\psi}{\rho}+\chi\left(\psi^{\prime}\cos\psi+\frac{\sin\psi}{\rho}\right)$
$+\kappa\left[2\psi^{\prime\prime\prime}\cos^{2}\psi-8\psi^{\prime\prime}\psi^{\prime}\cos^{2}\psi\sin\psi+\frac{4\psi^{\prime\prime}\cos^{3}\psi}{\rho}+2\left(\psi^{\prime}\right)^{3}\cos\psi\sin^{2}\psi\right]$
$+\kappa\left[-\cos^{3}\psi\left(\psi^{\prime}\right)^{3}-\frac{7\left(\psi^{\prime}\right)^{2}\cos^{2}\psi\sin\psi}{\rho}+\frac{\psi^{\prime}\sin^{2}\psi\cos\psi}{\rho^{2}}+\frac{2\cos^{2}\psi\sin\psi}{\rho^{3}}\right]$
$+\kappa\left[\frac{\sin^{3}\psi}{\rho^{3}}-\frac{2\psi^{\prime}\cos^{3}\psi}{\rho^{2}}+\frac{\psi^{\prime}\sin^{2}\psi\cos\psi}{\rho^{2}}\right],$
$A_{\beta}$ contains the terms related to $\beta^{(1)}_{1},\beta^{(1)}_{2},$ and $\beta_{\kappa}$
$A_{\beta}$ $=\beta_{1}^{(1)}\left(\frac{\omega^{4}}{16}\rho^{2}\cos^{2}\psi\right)+\beta_{2}^{(2)}\left(\frac{\omega^{4}}{8}\rho\cos^{2}\psi\sin\psi\right)$ (145)
$-\beta_{\kappa}\left(\frac{\omega^{4}}{8}\rho^{2}\cos^{2}\psi\right)\left(\psi^{\prime}\cos\psi+\frac{\sin\psi}{\rho}\right),$
$N^{(I)}_{i}$ contains terms associate to $n^{(I)}_{i}$
$N^{(1)}_{1}$ $=\frac{\omega^{2}}{8}n^{(1)}_{1}\left(3+\cos\left(2\psi\right)\right)\,,$ (146)
$N^{(2)}_{2}$ $=\frac{\omega^{2}}{2}n_{2}^{(2)}\cos^{2}\psi\left(\psi^{\prime}\cos\psi+\frac{\sin\psi}{\rho}\right),$ (147)
and $N_{\kappa}$ contains terms associated to $n_{\kappa}$
$N_{\kappa}$ $=-\frac{\omega^{2}}{4\rho}n_{\kappa}\sin^{3}\psi-2\omega^{2}n_{\kappa}\psi^{\prime}\cos^{3}\psi+\frac{5}{4}\omega^{2}n_{\kappa}\rho\left(\psi^{\prime}\right)^{2}\cos^{2}\psi\sin\psi-\omega^{2}n_{\kappa}\rho\psi^{\prime\prime}\cos^{3}\psi.$ (148)
In general, we will need the equations of state for $\{\chi,a^{(1)}_{1},\kappa,a^{(2)}_{2}\}$ to even attemp to solve this equation. The solution at $\omega=0$ and assuming constant transport coefficients is the well known biconcave discoid solutionCanham1970 ; Naito1996 ; Armas:2020
$\sin\psi$ $=\left(\frac{a_{1}^{(1)}}{2\kappa}\right)\rho\left[\ln\rho+b\right]\,,$ (149)
$\chi$ $=\kappa\left(\frac{a^{(1)}_{1}}{2\kappa}\right)^{2}.$ (150)
VI Conclusions
We have successfully used the formalism presented in Armas:2020 to construct the more general equilibrium partition function and, consequently, constitutive relations for NC membranes embedded in flat NC spacetime with both finite temperature and chemical potential. The equilibrium constraints imply that the temperature must be constant and the chemical potential must be related to the acceleration. The shape of the membrane in its equilibrium state will be determined by the generalized Young-Laplace equation (68). We wrote down this equation explicitely for a particular set of constitutive relations in the presence of a constant vorticity. We will leave the full study of solutions to this equation for future work. We will like to point out that in Anand2015 a relation between linear chemo-elasticity and linear poroelasticity was established, this means the presence of a mass chemical potential can be interpreted as having an underlying effective porous medium. This will allow us to eventually extend the range of applications of the resulting work in the context of finite temperature poroelastic membranes at hydrostatic equilibrium.
Acknowledgement
This work was supported by Universidad Nacional Autónoma de México through UNAM-DGAPA’s posdoctoral fellowship program ”Programa de Becas Posdoctorales en la UNAM.”
Appendix A
In here we show the variation of each of the terms appearing on the equilibrium partition function
A.1 Variation of $\chi$
The variation coming from the $\chi(T,\mu)$ term is explicitly given by
$\delta\mathcal{F}_{\chi}$ $=\int d^{d-1}\sigma\hat{e}\left[\frac{1}{2}\left(\chi h^{ab}+nu^{a}u^{b}\right)\delta\bar{h}_{ab}-\left(\chi\hat{\upsilon}^{c}+\left(\epsilon+\chi+\frac{n}{2}\bar{u}^{2}\right)u^{c}\right)\delta\tau_{c}\right].$ (151)
A.2 Variations of order one scalars
The variation of $F_{(1)}=K_{ab}h^{ab}$ is
$\delta F_{(1)}$ $=-\left(K_{ab}h^{ac}h^{bd}\right)\delta h_{cd}+\left(2\hat{\upsilon}^{a}h^{cb}K_{ab}\right)\delta\tau_{c}+\left(h^{cd}\right)\delta K_{cd},$ (152)
contributing to a variation of the equilibrium partition function given by
$\delta\mathcal{F}_{F_{(1)}}$ $=\int d^{d-1}\sigma\hat{e}\left[\frac{1}{2}\left(\xi^{(1)cd}_{1}F_{(1)}-2a_{1}^{(1)}h^{ac}h^{bd}K_{ab}\right)\delta\bar{h}_{cd}\right]$ (153)
$+\int d^{d-1}\sigma\hat{e}\left[\left(2a_{1}^{(1)}\hat{\upsilon}^{a}h^{bc}K_{ab}-\zeta^{(1)c}_{1}F_{(1)}\right)\delta\tau_{c}+\left(a_{1}^{(1)}h^{cd}\right)\delta K_{cd}\right].$
The variation of $F_{(2)}=K_{ab}u^{a}u^{b}$ is
$\delta F_{(2)}$ $=-\left(2K_{ab}u^{a}u^{b}u^{c}\right)\delta\tau_{c}+\left(u^{c}u^{d}\right)\delta K_{cd},$ (154)
contributing to a variation of the equilibrium partition function given by
$\delta\mathcal{F}_{F_{(2)}}$ $=\int d^{d-1}\sigma\hat{e}\left[\frac{1}{2}\left(\xi^{(1)cd}_{2}F_{(2)}\right)\delta\bar{h}_{cd}-\left(\zeta^{(1)c}_{2}F_{(2)}+2a_{2}^{(1)}K_{ab}u^{a}u^{b}u^{c}\right)\delta\tau_{c}\right]$ (155)
$+\int d^{d-1}\sigma\hat{e}\left(a^{(1)}_{2}u^{c}u^{d}\right)\delta K_{cd}.$
A.3 Variations of order two scalars
The variation of $S_{(1)}=\left(K_{ab}h^{ab}\right)\left(K_{cd}h^{cd}\right)$ is
$\delta S_{(1)}=2F_{(1)}\delta F_{(1)},$ (156)
meaning it will contribute to the variation of the equilibrium partition function as
$\delta\mathcal{F}_{S_{(1)}}$ $=\int d^{d-1}\sigma\hat{e}\left[\frac{1}{2}\left(\xi^{(2)cd}_{1}S_{(1)}-4a_{1}^{(2)}F_{(1)}h^{ac}h^{bd}K_{ab}\right)\delta\bar{h}_{cd}\right]$ (157)
$+\int d^{d-1}\sigma\hat{e}\left[\left(4a_{1}^{(2)}F_{(1)}\hat{\upsilon}^{a}h^{bc}K_{ab}-\zeta^{(2)c}_{1}S_{(1)}\right)\delta\tau_{c}+\left(2a_{1}^{(2)}F_{(1)}h^{cd}\right)\delta K_{cd}\right].$
The variation of $S_{(2)}=K_{ab}K_{cd}h^{ac}h^{bd}$ is
$\delta S_{(2)}$ $=\left(-2h^{ac}h^{bd}h^{rs}K_{ar}K_{bs}\right)\delta\bar{h}_{cd}+\left(4\hat{\upsilon}^{a}h^{bc}h^{dr}K_{ad}K_{br}\right)\delta\tau_{c}+\left(2h^{ac}h^{bd}K_{ab}\right)\delta K_{cd},$ (158)
and it will contribute to the variation of the equilibrium partition function as
$\delta\mathcal{F}_{S_{(2)}}$ $=\int d^{d-1}\sigma\hat{e}\left[\frac{1}{2}\left(\xi_{2}^{(2)cd}S_{(2)}-4a_{2}^{(2)}h^{ac}h^{bd}h^{rs}K_{ar}K_{bs}\right)\delta\bar{h}_{cd}\right]$ (159)
$+\int d^{d-1}\sigma\hat{e}\left[\left(4a_{2}^{(2)}\hat{\upsilon}^{a}h^{bc}h^{dr}K_{ad}K_{br}-\zeta_{2}^{(2)c}S_{(2)}\right)\delta\tau_{c}+\left(2a_{2}^{(2)}h^{ac}h^{bd}K_{ab}\right)\delta K_{cd}\right].$
The variation of $S_{(3)}=K_{ab}K_{cd}u^{a}h^{bc}u^{d}$ is
$\delta S_{(3)}$ $=\left(-h^{cr}h^{ds}K_{ar}K_{bs}u^{a}u^{b}\right)\delta h_{cd}+\left(2\hat{\upsilon}^{a}h^{cr}K_{ab}K_{dr}u^{b}u^{d}-2u^{c}h^{dr}K_{ad}K_{br}u^{a}u^{b}\right)\delta\tau_{c}$ (160)
$+\left(2u^{(c}h^{d)b}u^{a}K_{ab}\right)\delta K_{cd},$
and it will contribute to the variation of the equilibrium partition function as
$\delta\mathcal{F}_{S_{(3)}}$ $=\int d^{d-1}\sigma\hat{e}\left[\frac{1}{2}\left(\xi^{(2)cd}_{3}S_{(3)}-2a_{3}^{(2)}h^{cr}h^{ds}K_{ar}K_{bs}u^{a}u^{b}\right)\delta\bar{h}_{cd}\right]$ (161)
$+\int d^{d-1}\sigma\hat{e}\left[\left(2a_{3}^{(2)}\hat{\upsilon}^{a}h^{cr}K_{ab}K_{dr}u^{b}u^{d}-2a_{3}^{(2)}u^{c}h^{dr}K_{ad}K_{br}u^{a}u^{b}-\zeta^{(2)c}_{3}S_{(3)}\right)\delta\tau_{c}\right]$
$+\int d^{d-1}\sigma\hat{e}\left[\left(2a_{3}^{(2)}u^{(c}h^{d)b}u^{a}K_{ab}\right)\delta K_{cd}\right].$
The variation $S_{(4)}=\left(K_{ab}u^{a}u^{b}\right)\left(K_{cd}u^{c}u^{d}\right)$ is
$\delta S_{(4)}$ $=2F_{(2)}\delta F_{(2)},$ (162)
meaning it will contribute to the variation of the equilibrium partition function as
$\delta\mathcal{F}_{S_{(4)}}$ $=\int d^{d-1}\sigma\hat{e}\left[\frac{1}{2}\left(\xi^{(2)cd}_{4}S_{(4)}\right)\delta\bar{h}_{cd}-\left(\zeta_{4}^{(2)c}S_{(4)}+4a_{4}^{(2)}F_{(2)}K_{ab}u^{a}u^{b}u^{c}\right)\delta\tau_{c}\right]$ (163)
$+\int d^{d-1}\sigma\hat{e}\left(2F_{(2)}a_{4}^{(2)}u^{c}u^{d}\right)\delta K_{cd}.$
The variation of $S_{(5)}=a^{c}a^{d}\bar{h}_{cd}$ is
$\delta S_{(5)}=2\delta a^{c}a^{d}\bar{h}_{cd}+a^{c}a^{d}\delta\bar{h}_{cd},$ (164)
contributing to the variation of the equilibrium partition function as
$\delta\mathcal{F}_{S_{(5)}}$ $=\frac{a^{(2)}_{5}}{2}\int d^{d-1}\sigma\hat{e}\left[\frac{\xi_{5}^{(2)cd}S_{(5)}}{a^{(2)}_{5}}+2a^{c}a^{d}+4a_{f}h^{cf}h^{dg}D_{d}\mu+2h^{fg}D_{f}a_{g}u^{c}u^{d}\right]\delta\bar{h}_{cd}$ (165)
$-\int d^{d-1}\sigma\hat{e}\left[\left(\zeta_{5}^{(2)c}S_{(5)}+4a^{(2)}_{5}\hat{\upsilon}^{(f}h^{g)c}a_{f}D_{g}\mu+2a^{(2)}_{5}u^{c}\left(\mu+\frac{1}{2}\bar{u}^{2}\right)h^{fg}D_{f}D_{g}\mu\right)\delta\tau_{c}\right].$
The variation of $S_{(6)}=K_{cd}u^{c}a^{d}$ is
$\delta S_{(6)}=\left(a^{c}u^{d}\right)\delta K_{cd}-\left(K_{fg}a^{a}u^{g}u^{c}\right)\delta\tau_{c}+K_{cd}u^{c}\delta a^{d},$ (166)
contributing to the variation of the equilibrium partition function as
$\delta\mathcal{F}_{S_{(6)}}$ $=\int\frac{d^{d-1}\sigma\hat{e}}{2}\left[\xi^{(2)cd}_{6}S_{(6)}+a^{(2)}_{6}\left(2K_{ab}u^{a}h^{bc}h^{fd}D_{f}\mu+h^{bf}D_{f}\left(K_{ab}u^{a}\right)u^{c}u^{d}\right)\right]\delta\bar{h}_{cd}$ (167)
$-\int d^{d-1}\sigma\hat{e}\left(\zeta^{(2)c}_{6}S_{(6)}+a^{(2)}_{(6)}K_{fg}a^{f}u^{g}u^{c}+2a^{(2)}_{6}K_{ab}u^{a}\hat{\upsilon}^{(b}h^{f)c}D_{f}\mu\right)\delta\tau_{c}$
$+\int d^{d-1}\sigma\hat{e}\left[-h^{bf}D_{f}\left(a^{(2)}_{6}K_{ab}u^{a}\right)\left(\mu+\frac{1}{2}\bar{u}^{2}\right)u^{a}\delta\tau_{a}+\left(a^{(2)}_{6}a^{c}u^{d}\right)\delta K_{cd}\right].$
The variation of $S_{(7)}=\Omega^{ab}\Omega^{cd}\bar{h}_{ac}\bar{h}_{bd}$ is
$\delta S_{(7)}$ $=2\bar{h}_{ab}\Omega^{ac}\Omega^{bd}\delta\bar{h}_{cd}+2\bar{h}_{ac}\bar{h}_{bd}\Omega^{ab}\delta\Omega^{cd}$ (168)
$=\left[\bar{u}_{a}\bar{h}_{bc}\Omega^{ab}-D_{c}\left(\mu+\frac{1}{2}\bar{u}^{2}\right)\right]\Omega^{cd}\delta\tau_{d}-\left(\Omega^{af}\Omega^{bg}\bar{h}_{ab}\bar{h}_{fg}\right)u^{c}\delta\tau_{c}$ (169)
$+\frac{1}{2}\delta\bar{h}_{cd}\Omega^{cf}\Omega^{dg}\bar{h}_{fg}-\bar{h}_{bc}\Omega^{cf}u^{b}u^{d}D_{f}\delta\tau_{d}+\Omega^{cd}u^{b}D_{c}\delta\bar{h}_{bd},$
contributing to the variation of the equilibrium partition function as
$\delta\mathcal{F}_{S_{(7)}}$ $=\int\frac{d^{d-1}\sigma\hat{e}}{2}\left[\xi^{(2)cd}_{7}S_{(7)}+2a^{(2)}_{7}\Omega^{cf}\Omega^{dg}\bar{h}_{fg}-2D_{b}\left(a^{(2)}_{(7)}\Omega^{bd}u^{c}\right)\right]\delta\bar{h}_{cd}$ (170)
$-\int d^{d-1}\sigma\hat{e}\left[\zeta^{(2)d}_{7}S_{(7)}-a^{(2)}_{(7)}\bar{u}_{a}\bar{h}_{bc}\Omega^{ab}\Omega^{cd}+a^{(2)}_{7}D_{c}\left(\mu+\frac{1}{2}\bar{u}^{2}\right)\Omega^{cd}\right]\delta\tau_{d}$
$+\int d^{d-1}\sigma\hat{e}\left[D_{f}\left(a^{(2)}_{(7)}\bar{h}_{bc}\Omega^{cf}u^{b}\right)-a^{(2)}_{(7)}\left(\Omega^{af}\Omega^{bg}\bar{h}_{ab}\bar{h}_{fg}\right)\right]u^{c}\delta\tau_{c}.$